# Tree balance in phylogenetic models

### v3

@@ 2 Preliminaries
A very foundational tree model is the Yule model [61], which is a pure birth model, where all species have the same constant rate of speciation at all times. To obtain a tree shape under the Yule model using a forward process, starting with a single leaf, a leaf is chosen uniformly at random and replaced by a cherry (see Figure 1 on the left). Finally, a phylogenetic tree under the Yule model is obtained by assigning leaf labels uniformly at random to a shape generated by this process. In particular, this means that for the generated tree shape $T\in\mathcal{BT}^{\ast}_{n}$, each of the possible $\frac{n!}{2^{s(T)}}$ phylogenetic trees is equally likely to be chosen. In practice, this can for instance be achieved by choosing, for $i=1\ldots,n$, one leaf at a time uniformly at random from the set of leaves that have not been labeled yet and assigning it label $i$ [^2] Note that this procedure leads to a multiset of phylogenetic trees in which precisely $2^{s(T)}$ many (isomorphic) versions of each phylogenetic tree are contained (due to the symmetry nodes). Thus, when drawing one phylogenetic tree uniformly at random from this multiset, every phylogenetic tree has the same probability. The Yule model plays a central role in phylogenetics and has been used for many decades. It is known under a multitude of names (Yule model, equal-rates-Markov model (ERM), random branching model, Markovian dichotomous branching model, Yule-Harding model, Yule-Harding-Kingman model (YHK), or simply Markovian model) of which we will only use Yule model in the remainder of this manuscript. While it can be described under various approaches, the probabilities to obtain a specific (phylogenetic) tree can be explicitly stated as follows (see, e.g., [62, Proposition 1] or [63, Proposition 3.2]):
@@
In particular the formula for $P_{Y,n}(\mathcal{T})$ has been extensively studied in the literature and connections to other fields have been established (see, e.g., [64, 65, 66]).
Figure 1: Left: Illustration of the Yule model using the forward process (“tree growing”). At each step from $n$ to $n+1$, a leaf is chosen uniformly at random and replaced by a cherry. Thus, under the Yule model, the caterpillar tree on four leaves, occurs with probability $P_{Y,4}(T^{cat}_{4})=\frac{2}{3}$, whereas the fully balanced tree of height two occurs with probability $P_{Y,4}(T^{fb}_{2})=\frac{1}{3}$. Note that only considering $n=4$ leaves might lead to the wrong impression that the Yule model favors highly imbalanced trees. In fact, the caterpillar tree is actually getting less and less likely the more leaves are involved (as only the choice of any of the two leaves of the unique cherry in the caterpillar on $n$ leaves can generate a caterpillar on $n+1$ leaves). In fact, as the figure shows, already for $n=5$, the caterpillar is not the most likely tree anymore. Moreover, note that there are 12 phylogenetic trees of caterpillar shape and 3 phylogenetic trees of fully balanced shape. Thus, under the Yule model, the probability to observe a specific phylogenetic tree $\mathcal{T}$ of caterpillar shape is $P_{Y,4}(\mathcal{T})=\frac{2}{3}\cdot\frac{1}{12}=\frac{1}{18}$, and the probability to observe a specific phylogenetic tree $\mathcal{T}$ of fully balanced shape is $P_{Y,4}(\mathcal{T})=\frac{1}{3}\cdot\frac{1}{3}=\frac{1}{9}$.
@@
Aldous’ $\beta$-splitting model, $\beta\geq-2$ [67]
@@
Other
$\lambda_{i}$
Description and examples
Yule model [61, 16] $(\lambda_{0},\lambda_{0})$ $\lambda_{0}$ Pure birth process, constant and equal rates.
@@
Unequal fert. inheritance (IF-diff), simple biased speciation, $\zeta\geq 1$ [22, p. 1177] $\left(\frac{2\zeta}{\zeta+1}\lambda_{p},\frac{2}{\zeta+1}\lambda_{p}\right)$ $\lambda_{i}$ Factor $\zeta$ has long-time and exponential effect (opposite effect on the two children). Proportion $\lambda_{c_{2}}/\lambda_{c_{1}}=1/\zeta$ with $\lambda_{c_{1}}+\lambda_{c_{2}}=2\lambda_{p}$. Equals Yule for $\zeta=1$.
Biased speciation, $\zeta\in[0,1]$, [26, p. 149] $(\zeta\lambda_{p},(1-\zeta)\lambda_{p})$ $\lambda_{i}$ The parent’s rate is divided between its two children according to the given ratio $\zeta$. Note that $\zeta\in[0,0.5]$ yields the same model.
@@
$\left(10^{\log(x_{c_{1}})+\varepsilon_{3}}\right.$,
$\quad\left.10^{\log(x_{c_{2}})+\varepsilon_{4}}\right)$
@@
$x_{i}$ Inherited trait influences speciation rate. With $\varepsilon_{1,2}\sim N(0,\sigma^{2}_{x})$ and $\varepsilon_{3,4}\sim N(0,\sigma^{2}_{\lambda})$. Trait values $<0$ are set to a small value $>0$. Note that $\log=\log_{10}$. There is also a bounded version of this model with an upper limit for the trait values, e.g., $x_{i}\in(0,20)$ symmetrical around $x_{0}=10$ ($x_{0}=100$ in [31]). Equals Yule for $\sigma_{\lambda}=\sigma_{x}=0$.
@@
$\left(10^{\log(x_{c_{1}})+\varepsilon_{3}}\right.$,
$\quad\left.10^{\log(x_{c_{2}})+\varepsilon_{4}}\right)$
@@
$\left(10^{\log(x_{p})+\varepsilon_{1},}\right.$
$\left.10^{\log(x_{p})+\varepsilon_{2}}\right)$
$x_{i}$ Inherited trait (with proportional changes) influences speciation rate. With $\varepsilon_{1,2}\sim N(0,\sigma^{2}_{x})$ and $\varepsilon_{3,4}\sim N(0,\sigma^{2}_{\lambda})$. Note that $\log=\log_{10}$. Equals Yule for $\sigma_{\lambda}=\sigma_{x}=0$.
Table 5: List of discrete-time tree growing models with extinction. These models act at the current leaves and modify their speciation and extinction rates. The process starts with a single leaf (the root) with speciation rate $\lambda_{0}$ and extinction rate $\mu_{0}$. At each time step one of the currently existing leaves is chosen for a speciation or extinction event. In case of speciation, it is split into two leaves, $c_{1}$ and $c_{2}$. We denote the resulting rates of $c_{1}$ and $c_{2}$ as $(\lambda_{c_{1}},\lambda_{c_{2}})$ and $(\mu_{c_{1}},\mu_{c_{2}})$, respectively. $\lambda_{e}$ and $\mu_{e}$ denote the rates of the leaf chosen for an extinction event.
@@
(Simple/constant-rate) birth-death model, $\lambda_{0}>0$, $\mu_{0}\geq 0$ [61, 2]
@@
Alternative birth-death model, $\lambda_{0}>0$, $\mu_{0}\geq 0$
@@
Density-dependent model [68]
@@
SSE models state-dependent speciation and extinction models (see, e.g., [69, 70, 71, 72, 73, 74]) Family of models extending birth-death models to account for speciation, extinction, and trait evolution. Each species is associated with a specific trait/state (observed or hidden), and can transition to another state, or speciate (resp. go extinct) with rates depending on its current state.
Table 6: List of tree growing models that do not use rates.
Model Description
Ford’s $\alpha$-model, $\alpha\in[0,1]$ [49, 75] Start with a cherry and add an additional temporary root edge $(\rho_{temp},\rho)$. Then, in each step – let $m$ be the current leaf number – choose any edge, where each of the $m-1$ inner edges has probability $p=\frac{\alpha}{m-\alpha}$ and each of the $m$ pendant edges has probability $p=\frac{1-\alpha}{m-\alpha}$, subdivide it, and add a new pendant edge with a leaf in each step. Do this until reaching $n$ leaves. Last, delete the edge incident to $\rho_{temp}$. Equals Yule for $\alpha=0$, PDA for $\alpha=1/2$, comb for $\alpha=1$.
Uniform model, proportional-to-distinguishable-arrangements (or -types) model (PDA) [17] Start with an unrooted tree with 3 leaves. Then, in each step, choose any edge, subdivide it, and add a new pendant edge with a leaf until reaching $n$ leaves. Last, choose an edge to insert a root [76, p. 8].
Table 7: List of distribution-based tree models. These models generate trees according to some probability distribution.
Model Probabilities Description
Uniform model, proportional-to-distinguishable-arrangements (or -types) model (PDA) [17]
@@
Uniform model, equiprobable-types-model (ETM)
@@ 3 Methods: How to measure the power of balance indices
While there are various ways of comparing tree balance indices, in the present manuscript, we focus on comparing their power to distinguish alternative phylogenetic models from a null model $P^{0}_{n}$. As most previous studies (e.g., [32, 22, 33, 31, 26, 34, 35, 36]) we choose the Yule model as the null model for the main paper (we present some additional results using other null models in the supplementary material [40]). The Yule model also belongs to several parameterized tree model families, which allows us to observe the power of TSS as we gradually move away from the Yule model by changing the respective parameter(s). The null hypothesis $H_{0}$ is that the trees have been constructed under the null model.
@@
    Compute or approximate null distribution of the TSS values: If $P^{0}_{n}$ is either the Yule model, the PDA model, or the ETM, we can compute the exact distribution (function) of $t$ under $P^{0}_{n}$ for smaller $n$ (our package poweRbal provides this option for $n\leq 20$) by generating all trees $T\in\mathcal{BT}^{\ast}_{n}$ with $n$ leaves, calculating their TSS values and using the probabilities $P_{Y,n}(T)$ given in Equation (1) for the Yule model, and $P_{\textup{PDA},n}(T)$ and $P_{\textup{ETM},n}(T)$ given in Table 7 for the PDA and ETM models.
    For larger $n$ (or if $P^{0}_{n}$ is not Yule, PDA, or ETM) we approximate the distribution (function) by generating $N_{d}=10^{4}$ to $10^{5}$ trees under the null model $P^{0}_{n}$ and computing their TSS values (see Figure 2).
@@
    for a lower and upper bound $c_{1},c_{2}\in\mathbb{R}$. $c_{1}$ is chosen as the largest TSS value of step (1) with $P(t(T)<c_{1}|H_{0})\leq\frac{\alpha}{2}=2.5\%$ and $c_{2}$ as the smallest TSS value with $P(t(T)>c_{2}|H_{0})\leq\frac{\alpha}{2}=2.5\%$. Furthermore, we do a small-sample correction [77, Equation  21] to ensure that $P(\Phi(t(T))=1|H_{0})$ is as close to $\alpha$ as possible (and not lower, which would make the test too conservative). For this purpose, we compute the probabilities $p_{1}=\frac{\alpha/2-P(t(T)<c_{1}|H_{0})}{P(t(T)=c_{1}|H_{0})}$ and $p_{2}=\frac{\alpha/2-P(t(T)>c_{2}|H_{0})}{P(t(T)=c_{2}|H_{0})}$ to reject the null hypothesis if the TSS value equals the lower or upper bound.
@@
    Test trees under alternative model: Construct $N_{a}=10^{3}$ many trees $(T_{i}^{a})_{i=1,...,N_{a}}$ under $P^{a}_{n}$. For each tree $T_{i}^{a}$ compute $t(T_{i}^{a})$ and $\Phi(t(T_{i}^{a}))$.
@@
Figure 2: Histograms of the exact distributions of (a) the Furnas rank and (b) the Colless index under the Yule model, the PDA model, and the ETM for $n=16$. The Furnas rank is a balance index, i.e., trees with higher values can be considered as more balanced, and the Colless index is an imbalance index, i.e., trees with higher values can be considered as more imbalanced. The uniform distribution of the ETM on $\mathcal{BT}^{\ast}_{n}$ is directly visible as the Furnas rank is also a tree enumeration method, which assigns all trees in $\mathcal{BT}^{\ast}_{n}$ a unique number in $\{1,...,|\mathcal{BT}^{\ast}_{n}|\}$ (here $|\mathcal{BT}^{\ast}_{16}|=10,905$). The dotted lines indicate the 0.025- and the 0.975-quantile under the Yule model. In Panel (c), histograms of the distributions of the Colless index under the Aldous’ $\beta$ splitting model for several choices of $\beta$ (2,000 sampled trees for each $\beta$ in this case) for $n=30$ are shown with overlaid density curves, estimated using Gaussian kernel density estimation.
@@
While the cherry index is not a tree balance index [14], it is included in our study as it has been used in similar contexts already [26, 43]. The complete list of TSS can be found in the legends of the graphics.
Furthermore, we use (representatives of) all discrete-time tree models $P_{n}$ listed in Section 3 in the supplementary material [40]. The only exceptions are the most simplistic models that only produce a single tree shape, e.g., the caterpillar or maximally balanced tree. The comparisons are done for $n=30$ and $100$ to cover some range of tree sizes.

### v4

@@ I-22 Cherry index [62] number of cherries in $T$ $ChI(T)\coloneqq c(T)$
A very foundational tree model is the Yule model [63], which is a pure birth model, where all species have the same constant rate of speciation at all times. To obtain a tree shape, i.e., an unlabeled (rooted binary) tree, under the Yule model using a forward process, starting with a single leaf, a leaf is chosen uniformly at random and replaced by a cherry (see Figure 1 on the left). Finally, a phylogenetic tree, i.e., a leaf-labeled tree, under the Yule model is obtained by assigning leaf labels uniformly at random to a shape generated by this process. In particular, this means that for the generated (unlabeled) tree shape $T\in\mathcal{BT}^{\ast}_{n}$, each of the possible $\frac{n!}{2^{s(T)}}$ (leaf-labeled) phylogenetic trees is equally likely to be chosen. In practice, this can for instance be achieved by choosing, for $i=1\ldots,n$, one leaf at a time uniformly at random from the set of leaves that have not been labeled yet and assigning it label $i$ [^1] Note that this procedure leads to a multiset of phylogenetic trees in which precisely $2^{s(T)}$ many (isomorphic) versions of each phylogenetic tree are contained (due to the symmetry nodes). Thus, when drawing one phylogenetic tree uniformly at random from this multiset, every phylogenetic tree has the same probability. The Yule model plays a central role in phylogenetics and has been used for many decades. It is known under a multitude of names (Yule model, equal-rates-Markov model (ERM), random branching model, Markovian dichotomous branching model, Yule-Harding model, Yule-Harding-Kingman model (YHK), or simply Markovian model) of which we will only use Yule model in the remainder of this manuscript. While it can be described under various approaches, the probabilities of obtaining a specific (phylogenetic) tree can be explicitly stated as follows (see, e.g., [64, Proposition 1] or [65, Proposition 3.2]):
@@
In particular, the formula for $P_{Y,n}(\mathcal{T})$ has been extensively studied in the literature and connections to other fields have been established (see, e.g., [66, 67, 68]).
Figure 1: Left: Illustration of the Yule model using the forward process (“tree-growing”). At each step from $n$ to $n+1$, a leaf is chosen uniformly at random and replaced by a cherry. Thus, under the Yule model, the caterpillar tree on four leaves occurs with probability $P_{Y,4}(T^{cat}_{4})=\frac{2}{3}$, whereas the fully balanced tree of height two occurs with probability $P_{Y,4}(T^{fb}_{2})=\frac{1}{3}$. Note that only considering $n=4$ leaves might lead to the wrong impression that the Yule model favors highly imbalanced trees. In fact, the caterpillar tree is actually getting less and less likely the more leaves are involved (as only the choice of any of the two leaves of the unique cherry in the caterpillar on $n$ leaves can generate a caterpillar on $n+1$ leaves). In fact, as the figure shows, already for $n=5$, the caterpillar is not the most likely tree anymore. Moreover, note that there are 12 phylogenetic trees of caterpillar shape and 3 phylogenetic trees of fully balanced shape. Thus, under the Yule model, the probability to observe a specific phylogenetic tree $\mathcal{T}$ of caterpillar shape is $P_{Y,4}(\mathcal{T})=\frac{2}{3}\cdot\frac{1}{12}=\frac{1}{18}$, and the probability to observe a specific phylogenetic tree $\mathcal{T}$ of fully balanced shape is $P_{Y,4}(\mathcal{T})=\frac{1}{3}\cdot\frac{1}{3}=\frac{1}{9}$.
@@
M-1 ALDOUS’ $\beta$-splitting model, $\beta\geq-2$ [69]
@@
Other $\lambda_{i}$ Description and examples
M-0 YULE model [63, 16] $(\lambda_{0},\lambda_{0})$ $\lambda_{0}$ Pure birth process, constant and equal rates.
@@
M-4 Unequal fert. inheritance (IF-DIFF), simple biased speciation, $\zeta\geq 1$ [22, p. 1177] $\left(\frac{2\zeta}{\zeta+1}\lambda_{p},\frac{2}{\zeta+1}\lambda_{p}\right)$ $\lambda_{i}$ Factor $\zeta$ has long-time and exponential effect (opposite effect on the two children). Proportion $\lambda_{c_{2}}/\lambda_{c_{1}}=1/\zeta$ with $\lambda_{c_{1}}+\lambda_{c_{2}}=2\lambda_{p}$. Equals Yule for $\zeta=1$.
M-5 BIASED SPECIATION, $\zeta\in[0,1]$, [26, p. 149] $(\zeta\lambda_{p},(1-\zeta)\lambda_{p})$ $\lambda_{i}$ The parent’s rate is divided between its two children according to the given ratio $\zeta$. Note that $\zeta\in[0,0.5]$ yields the same model.
@@
$\left(10^{\log_{10}(x_{c_{1}})+\varepsilon_{3}}\right.$,
$\quad\left.10^{\log_{10}(x_{c_{2}})+\varepsilon_{4}}\right)$
@@
$x_{i}$ Inherited trait influences speciation rate. With $\varepsilon_{1,2}\sim N(0,\sigma^{2}_{x})$ and $\varepsilon_{3,4}\sim N(0,\sigma^{2}_{\lambda})$. Trait values $<0$ are set to a small value $>0$. There is also a BOUNDED version of this model with an upper limit for the trait values, e.g., $x_{i}\in(0,20)$ symmetrical around $x_{0}=10$ ($x_{0}=100$ in [31]). Equals Yule for $\sigma_{\lambda}=\sigma_{x}=0$.
@@
$\left(10^{\log_{10}(x_{c_{1}})+\varepsilon_{3}}\right.$,
$\quad\left.10^{\log_{10}(x_{c_{2}})+\varepsilon_{4}}\right)$
@@
$\left(10^{\log_{10}(x_{p})+\varepsilon_{1},}\right.$
$\left.10^{\log_{10}(x_{p})+\varepsilon_{2}}\right)$
$x_{i}$ Inherited trait (with proportional changes) influences speciation rate. With $\varepsilon_{1,2}\sim N(0,\sigma^{2}_{x})$ and $\varepsilon_{3,4}\sim N(0,\sigma^{2}_{\lambda})$. Equals Yule for $\sigma_{\lambda}=\sigma_{x}=0$.
Table 5: List of discrete-time tree-growing models with extinction. These models act at the current leaves and modify their speciation and extinction rates. The process starts with a single leaf (the root) with speciation rate $\lambda_{0}$ and extinction rate $\mu_{0}$. At each time step one of the currently existing leaves is chosen for a speciation or extinction event. In case of speciation, it is split into two leaves, $c_{1}$ and $c_{2}$. We denote the resulting rates of $c_{1}$ and $c_{2}$ as $(\lambda_{c_{1}},\lambda_{c_{2}})$ and $(\mu_{c_{1}},\mu_{c_{2}})$, respectively, and list them in the column labeled “Children”. $\lambda_{i}$ and $\mu_{i}$ denote the rates of all other species and are given in the column labeled “Other”. $\lambda_{e}$ and $\mu_{e}$ denote the rates of the leaf chosen for an extinction event (Column “Extinct”).
ID
@@
M-10 (Simple/constant-rate) BIRTH-DEATH model, $\lambda_{0}>0$, $\mu_{0}\geq 0$ [63, 2]
@@
M-11 ALTERNATIVE BIRTH-DEATH model, $\lambda_{0}>0$, $\mu_{0}\geq 0$
@@
M-12 DENSITY-dependent model [70]
@@
M-13 SSE models state-dependent speciation and extinction models (see, e.g., [71, 72, 73, 74, 75, 76]) Family of models extending birth-death models to account for speciation, extinction, and trait evolution. Each species is associated with a specific trait/state (observed or hidden), and can transition to another state, or speciate (resp. go extinct) with rates depending on its current state.
Table 6: List of tree-growing models that do not use rates.
ID Model Description
M-14 FORD’S $\alpha$-model, $\alpha\in[0,1]$ [51, 77] Start with a cherry and add an additional temporary root edge $(\rho_{temp},\rho)$. Then, in each step – let $m$ be the current leaf number – choose any edge, where each of the $m-1$ inner edges has probability $p=\frac{\alpha}{m-\alpha}$ and each of the $m$ pendant edges has probability $p=\frac{1-\alpha}{m-\alpha}$, subdivide it, and add a new pendant edge with a leaf in each step. Do this until reaching $n$ leaves. Last, delete the edge incident to $\rho_{temp}$. Equals Yule for $\alpha=0$, PDA for $\alpha=1/2$, comb for $\alpha=1$.
M-15 Uniform model, proportional-to-distinguishable-arrangements (or -types) model (PDA) [17] Start with an unrooted tree with 3 leaves. Then, in each step, choose any edge, subdivide it, and add a new pendant edge with a leaf until reaching $n$ leaves. Last, choose an edge to insert a root [78, p. 8].
Table 7: List of distribution-based tree models. These models generate trees according to some probability distribution. The PDA model here is the same as in Table 6.
ID Model Probabilities Description
M-15 Uniform model, proportional-to-distinguishable-arrangements (or -types) model (PDA) [17]
@@
M-16 Uniform model, equiprobable-types-model (ETM)
@@ 3 Methods: How to measure the power of balance indices
While there are various ways of comparing tree balance indices, in the present manuscript, we focus on comparing their power to distinguish alternative phylogenetic models from a null model $P^{0}_{n}$. As most previous studies (e.g., [32, 22, 33, 31, 26, 34, 35, 36]) we choose the Yule model M-0 as the null model for the main paper (we present some additional results using other null models in the supplementary material [40]). The Yule model also belongs to several parameterized tree model families, which allows us to observe the power of TSS as we gradually move away from the Yule model by changing the respective parameter(s). The null hypothesis $H_{0}$ is that the trees have been constructed under the null model.
@@
    Compute or approximate null distribution of the TSS values: If $P^{0}_{n}$ is either the Yule model M-0, the PDA model M-15, or the ETM M-16, we can compute the exact distribution (function) of $t$ under $P^{0}_{n}$ for smaller $n$ (our package poweRbal provides this option for $n\leq 20$) by generating all trees $T\in\mathcal{BT}^{\ast}_{n}$ with $n$ leaves, calculating their TSS values and using the probabilities $P_{Y,n}(T)$ given in Equation (1) for the Yule model, and $P_{\textup{PDA},n}(T)$ and $P_{\textup{ETM},n}(T)$ given in Table 7 for the PDA and ETM models.
    For larger $n$ (or if $P^{0}_{n}$ is not Yule, PDA, or ETM) we approximate the distribution (function) by generating $N_{d}=10^{5}$ (except for the PDA and ETM where we use $N_{d}=10^{4}$ to reduce the overall computation time for the many choices of $n$; cf. Figure 3) trees under the null model $P^{0}_{n}$ and computing their TSS values (see Figure 2).
@@
    for a lower and upper bound $c_{1},c_{2}\in\mathbb{R}$. $c_{1}$ is chosen as the largest TSS value of step (1) with $P(t(T)<c_{1}|H_{0})\leq\frac{\alpha}{2}=2.5\%$ and $c_{2}$ as the smallest TSS value with $P(t(T)>c_{2}|H_{0})\leq\frac{\alpha}{2}=2.5\%$. Furthermore, we do a small-sample correction [79, Equation  21] to ensure that $P(\Phi(t(T))=1|H_{0})$ is as close to $\alpha$ as possible (and not lower, which would make the test too conservative). For this purpose, we compute the probabilities $p_{1}=\frac{\alpha/2-P(t(T)<c_{1}|H_{0})}{P(t(T)=c_{1}|H_{0})}$ and $p_{2}=\frac{\alpha/2-P(t(T)>c_{2}|H_{0})}{P(t(T)=c_{2}|H_{0})}$ to reject the null hypothesis if the TSS value equals the lower or upper bound.
@@
    Test trees under alternative model: Construct $N_{a}=10^{3}$ trees $(T_{i}^{a})_{i=1,...,N_{a}}$ under $P^{a}_{n}$. For each tree $T_{i}^{a}$ compute $t(T_{i}^{a})$ and $\Phi(t(T_{i}^{a}))$.
@@
Figure 2: Histograms of the exact distributions of (a) the Furnas rank I-3 and (b) the Colless index I-9 under the Yule model M-0, the PDA model M-15, and the ETM M-16 for $n=16$. The Furnas rank is a balance index, i.e., trees with higher values can be considered as more balanced, and the Colless index is an imbalance index, i.e., trees with higher values can be considered as more imbalanced. The uniform distribution of the ETM on $\mathcal{BT}^{\ast}_{n}$ is directly visible as the Furnas rank is also a tree enumeration method, which assigns all trees in $\mathcal{BT}^{\ast}_{n}$ a unique number in $\{1,...,|\mathcal{BT}^{\ast}_{n}|\}$ (here $|\mathcal{BT}^{\ast}_{16}|=10,905$). The dotted lines indicate the 0.025- and the 0.975-quantile under the Yule model. In Panel (c), histograms of the distributions of the Colless index under the Aldous’ $\beta$ splitting model M-1 for several choices of $\beta$ (2,000 sampled trees for each $\beta$ in this case) for $n=30$ are shown with overlaid density curves, estimated using Gaussian kernel density estimation.
@@
While the cherry index I-22 is not a tree balance index [14], it is included in our study as it has been used in similar contexts already [26, 43]. The complete list of included TSS can be found in the legends of the graphics.
Furthermore, we use (representatives of) all discrete-time tree models $P_{n}$ listed in Section 3 in the supplementary material [40]. The only exceptions are the most simplistic models that only produce a single tree shape, e.g., the caterpillar or maximally balanced tree (M-17 and M-18). The comparisons are done for $n=30$ and $100$ to cover some range of tree sizes.